OpenStudy (anonymous):

What is the value of the fourth term in a geometric sequence for which a1 = 10 and r = 0.5? a4 = 1.25 right?

OpenStudy (agent47):

a1 = 10 a2 = 10*0.5 = 5 a3 = 5*0.5 = 2.5 a4 = 2.5*0.5 = 1.25

OpenStudy (agent47):

yes, you are correct